# NaturalBench: Evaluating Vision-Language Models on Natural Adversarial Samples

### v1

@@ 3 Collecting NaturalBench
Figure 5: Performance of GPT-3.5, LLaVA-1.5, and GPT-4o on NaturalBench. We also split NaturalBench (the English subset) into equal-sized training and test sets, and report zero-shot (in blue) and finetuned (in green) results. We report group accuracy (G-Acc) (introduced in Section 4), which awards a point when all four (image, question) pairs are answered correctly. We highlight key results: (1) Blind GPT-3.5 fails to surpass random chance performance (red dotted line), regardless of finetuning. (2) LLaVA-1.5 improves by $9\%$ by finetuning on NaturalBench’s training images. (3) Even GPT-4o gains $10\%$ G-Acc through vision finetuning on NaturalBench. These findings confirm that NaturalBench is a more vision-centric benchmark, and a potentially useful dataset for improving already advanced VLMs.
@@ 5 Why is NaturalBench Challenging?
DeepSeek-VL-7B-Chat SigLIP-L SAM-B 45.8 86.6 49.9 81.8 19.4 54.8

### v2

@@ 3 Collecting NaturalBench
Figure 5: Performance of GPT-3.5, LLaVA-1.5, and GPT-4o on NaturalBench. We also split NaturalBench (the English subset) into equal-sized training and test sets, and report zero-shot (in blue) and finetuned (in green) results. We report group accuracy (G-Acc) (introduced in Section 4), which awards a point when all four (image, question) pairs are answered correctly. We highlight key results: (1) Blind GPT-3.5 fails to surpass random chance performance (red dotted line), regardless of finetuning. (2) LLaVA-1.5 improves by $9\%$ by finetuning on NaturalBench’s training images. (3) Even GPT-4o gains $10\%$ G-Acc through vision finetuning on NaturalBench; however, it falls far behind human performance (purple dotted line). These findings confirm that NaturalBench is a more vision-centric benchmark, and a potentially useful dataset for improving already advanced VLMs.
@@ 5 Why is NaturalBench Challenging?
DeepSeek-VL-7B-Chat SigLIP-L DeepSeek-LLM-7B 45.8 86.6 49.9 81.8 19.4 54.8
@@ Appendix D Debiasing Analysis
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] The main claims are reflected in the paper’s contributions and scope. See Section 3 to Section 6.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 5.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Section 5.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Section 3 for detailed benchmark collection pipeline. We have released the code in our project site.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See supplement.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [N/A]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See supplement.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [N/A]
    2. (b)
      Did you mention the license of the assets? [N/A]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [N/A]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [No] The assets used are all publicly sourced, and therefore explicit consent was not required.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [No] The data used are publicly available datasets that do not contain offensive content and with consent for personally identifiable information.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Section 3.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] We pay the participants above the minimum wage with an hourly pay.